Electrical non-conformities on instrumentation cables of the power converters will be closed by the end of the week.
The procedure for cleaning the DFBXH is currently being completed. In principle, the instrumentation wont have to be removed.
Planned activities
PIC-PC interface test - Wk 43
Room temperature vacuum and collimator works (wks 43-46), bake-out (wks 47 -48)
ATLAS-RP cable installation - Wks 48-49
TS/EL:
DFBXH Ineo cabling - To be done on Monday 22.
Installation of the removed water-cooled cable and RQ6 - Flohe Wk 44 - 45 Once the cleaning is completed.
NC DFBXH: The water cooled cable associated to the current lead #3 cannot be fixed to the lead on the left side. There is a mechanical interference between the water tube on the lug and a connector necessary for the routing of the protection signal from the intrumentation port. The connector cannot be moved to another position, it is therefore proposed to fix the water cooled cable on the right side of the current lead lug.
ElQA
TP4-A: Inner triplet L1 - as soon as possible since it will be the first time they are performed
TP4-A: Inner triplet R8 - not before the DFBX is repaired
Alarm test in the whole sector will be carried out on Wednesday 24.
General
Cooling water shut-down - Wks 42-44 - All the cryogeninc activities will be stopped.
A latest date for taking a decission about repairing or not the DFBAs has to be set for not jeopardizing the starting date of the flushing